OpenStudy (anonymous):

if the volume of a right circular cone is 324p cubic inches and the radius of the base is 9 inches long, find the height of the cone

OpenStudy (anonymous):

\[V = \dfrac 1 3 \pi r^2 h\]

OpenStudy (anonymous):

V = 324p cubic inches

OpenStudy (anonymous):

radius of the base is 9 inches

OpenStudy (anonymous):

\[h =\frac{ 3V }{ \pi r^2 }\]

OpenStudy (anonymous):

now you can plug in the values yourself